# FRIEDA: Benchmarking Multi-Step Cartographic Reasoning in Vision-Language Models

### v1

@@ 4.1 Experimental Setup
We evaluate 11 LVLMs with multi-image support on FRIEDA. For proprietary models, we test three models: Gemini-2.5-Pro (c:gemini2025gemini25pro), GPT-5-Think (c:OpenAI2025gpt5), and Claude-Sonnet-4 (c:anthropic2025sonnet). For open source models, we consider eight model families and evaluate the largest available model from each family: LLaVA-NeXT-110B (c:li2024llavanext), GLM4.5V-108B (c:vteam2025glm45v), InternVL3-78B (c:chen2024internvl3), LLaVA-OneVision-72B (c:li2024llavaov), Qwen2.5VL-72B (bai2025qwen25vl), InternVL3.5-38B (c:wang2025internvl35), [^2] We evaluate the 38B variant rather than the 241BA28B variant as the latter activates only 28B parameters during inference. We report the results for the 241BA28B setting in Appendix F.2. Ovis2-34B (c:lu2024ovis), and Ovis2.5-9B (c:lu2025ovis25).
To enforce determinism in open-source models, we set do_sample=False and temperature=0. For proprietary models, we use the default settings of each model with maximum reasoning enabled (e.g., reasoning=high for GPT-5-Think) and append the instruction “Do not use search” to turn off external retrieval. All models receive the same set of instructions that human annotators receive (Appendix B.4).
@@
Answers in FRIEDA fall into three categories: textual, distance, and direction. For textual answers, we employ an LLM-as-Judge (c:gu2025surveyllmasajudge) method, utilizing Mistral Small 3.1 (c:mistral2024) as the evaluator. [^3] Mistral is not the language backbone of any tested LVLM, thereby reducing potential bias c:panickssery2024preference. The full judge prompt appears in Appendix E.1. This setup handles minor variation (e.g., ‘Cypress Creek’ vs. ‘Cypress’) by matching semantics rather than identifying exact string equality. For distance-based answer, we report mean absolute percentage error (MAPE) and consider predictions within 20% error as correct, following c:lewis1982mape. For directional answers, we mark a response correct if it matches the target cardinal direction within one adjacent label (e.g., if the gold answer is North, accept North, North West, and North East), reflecting the perceptual nature of the labels. We validate the reliability of the evaluation method against manual annotations, achieving a Cohen’s $\kappa$ of 0.9028 across all judged questions, which supports its suitability for evaluation.
@@ 4.2 Evaluation Results
Figure 3 summarizes the overall performance, and Table 2 reports accuracy by spatial relation. As FRIEDA retains questions with at least 2/3 annotator agreeing on the gold answer, we report accuracy for two subsets: All-Agree, where all three annotators agreed, and Partial-Agree, where 2/3 annotators agreed. All-Agree items serve as an indirect indicator of questions that are easier and less ambiguous for the annotators under our task and instructions, whereas Partial-Agree items may be considered as intrinsically more difficult or ambiguous to answer correctly. We also report the Overall Accuracy, which aggregates both subsets. Even the strongest LVLM (Gemini-2.5-Pro) remains below 40% overall accuracy, well behind human performance at 84%. The best open source result (Ovis2.5-9B-Think) achieves 24% overall accuracy, underperforming proprietary systems and far below humans. We find no clear relationship between model size and performance, suggesting that training data, training objectives, and explicit reasoning mechanisms matter more than scale for cartographic reasoning.
@@
Overall
Border
Distance
Equal
Intersect
Orientation
Within
@@ 5 Analysis
To pinpoint where LVLMs fail, we analyze Gemini-2.5-Pro on the All-Agree subset (in total, 167 questions). This ensures that our analysis targets distinct model failures on questions that humans find straightforward. We assign each incorrect answer to a single primary error category. When multiple issues co-occur, we prioritize errors that occur earlier in the reasoning pipeline that propagate to downstream steps. The largest source of error involves the misinterpretation of legends (25.61%): cases where the model assigns colors or symbols to the wrong class. The remaining 23.78% is due to cross-map interpretation failures, which reflect difficulties in aligning the map scales and shared features across maps, and 16.46% is due to spatial-relation semantics error, which arises when the model mixes up spatial relations (e.g., labeling region B within A when it only touches A at the boundary). Map-element misunderstandings include mistakes with the map scale (9.76%; unit or ratio errors), map text (8.93%; selecting the wrong place or feature from labels), geometry or shape reference (3.66%; pointing to the wrong area on the map), and orientation (3.05%; ignoring a tilted compass). Finally, we observe generic VQA errors not specific to cartography, such as miscounting (6.71%), subject-object confusion (1.82%; referring ‘A relative to B’ as ‘B relative to A’), and hallucination (1.20%). For the top three error categories, we provide examples and rationales returned by the three proprietary models in Appendix F.1.
Figure 4: Per spatial relation accuracy (%) of human annotators and three proprietary LVLMs (Gemini-2.5-Pro, Claude-Sonnet-4, and GPT-5-Think) on FRIEDA-direct.
@@
Figure 4 reports per-spatial relation accuracy for human annotators and the three proprietary models. LVLM performance broadly tracks the human baseline: both are most accurate on orientation and struggle most with distance. On questions where an annotator answers incorrectly, LVLMs are also incorrect 84.53% of the time. While GPT-5-Think and Gemini-2.5-Pro achieve comparable overall accuracy, GPT-5-Think is stronger on tasks that require multi-map reasoning (Table 10), indicating better integration of evidence across maps. This is most evident in the equal relation questions, a multi-map exclusive task, where GPT-5-Think’s accuracy is nearly 13% higher compared to Gemini-2.5-Pro. Notably, Claude-Sonnet-4 is the strongest on distance questions, particularly those requiring interpretation of the map scale to compute exact distances.
@@
We observe a minimal difference in accuracy between the FRIEDA-direct (Figure 3 and Table 2) and FRIEDA-contextual (Appendix E.3). To verify that this is not an artifact of the accuracy metric, we directly compare the per-question performance of the eight open-source models under deterministic settings (i.e., do_sample=False and temperature=0). We observe 88.03% per-question performance agreement between the direct and contextual settings, indicating that contextual images (maps from the same document that are not required to answer the question) rarely affect the model’s prediction.
@@
Despite being the smallest model tested with FRIEDA, Ovis2.5-9B-Think attains strong results (4th overall and 1st among open source models). To identify what drives this performance, we further evaluate Ovis2.5-9B with explicit reasoning (i.e., Think) disabled (Table 3). The overall accuracy of Ovis2.5-9B remains above the open source average, indicating that model characteristics (e.g., architecture, training data) contribute to the model’s strong results. Enabling Think adds an additional 5% performance gain. To identify which question types benefit from explicit reasoning and whether it improves cartographic performance, we manually analyze the 60 questions that only the Think variant answers correctly. Reasoning helps mostly with cardinal-direction questions, where north faces the top of the image (48.33%), followed by multi-map alignment (23.33%). Additional improvements come from correctly reading map text (15%), interpreting the map scale (5%), associating legend with symbol (5%), and counting (3%). Together, these patterns suggest that explicit reasoning primarily strengthens orientation-related and multi-map questions, which are central to carographic reasoning, while yielding smaller gains in symbol and map scale interpretation. [^4] We further evaluate the association between performance and model size in Appendix F.2
@@ 6 Related Work
Recent benchmarks established baselines for LVLM reasoning over documents and designed graphics. In the document domain, DocVQA (c:matthew2021docvqa) and DocVXQA (c:souibgui2025docvxqa) introduce a large-scale question-answering (QA) dataset over real forms and reports, while DocoPilot (c:duan2025docopilot) extends evaluation to scientific articles, which involve embedded figures. For graphics, InfographicsVQA (c:matthew2022infovqa) and InfoChartQA (c:lin2025infochartqa) test reasoning over rich layouts and charts. In general, frontier LVLMs reveal competence at high-level patterns, such as trends and extrema, but struggle with precise value extraction and robustness. FRIEDA evaluates these shortcomings in a cartographic setting where layout, symbols, legends, scales, and compass orientation interact tightly to measure how well LVLMs integrate these signals to answer map-based questions.
@@
While recent map VQA benchmarks have expanded the scope of evaluation, they remain constrained to single-map tasks or specific domains. MapQA (c:chang2022mapqa) evaluates question answering on choropleth maps and shows that general VQA and ChartVQA systems underperform on maps. MapWise (c:mukhopadhyay2025mapwise) broadens the geographic coverage, while MapIQ (c:srivastava2025mapiq) extends the map type coverage to include cartograms and proportional-symbol maps. MapEval (c:dihan2025mapeval) assesses geospatial reasoning across various cities, and it reports consistent human-LVLMs performance gaps. Domain-specific efforts include PEACE (c:huang2025peace) for geology, CartoMark (c:zhou2024cartomark) for text extraction and recognition, and MapBench (c:xing2025mapbench) and ReasonMap (c:feng2025canmllmguidemehome) for navigation. However, these benchmarks rarely test cross-image reasoning on heterogeneous sources and often rely on a limited set of spatial relations. While ReMI (c:kazemi2025remi) explores the cross-image setting, the questions lack cartographic focus. We detail key differences between prior map VQA benchmarks and FRIEDA in Appendix G.
Spatial reasoning benchmarks, such as SpatialVLM (c:chen2024spatialvlm) and SpatialRGPT (c:cheng2024spatialrgpt), have advanced spatial perception and reasoning on natural images. In the geospatial domain, GeoChain (c:yerramilli2025geochain) enhances tasks like geolocalization by inducing step-by-step geographic reasoning. However, these works do not engage with symbolic conventions unique to maps (i.e., legends, scales, compasses, and map texts). In contrast, our benchmark closes this gap by evaluating multi-step cartographic reasoning over heterogeneous, real-document maps, which requires models to integrate evidence across multiple figures and align legends, scales, and orientation to infer key spatial relations (i.e., border, distance, equal, intersect, orientation, and within).
@@ 7 Conclusion
We present FRIEDA, a benchmark for evaluating multi-step cartographic reasoning across six spatial relations, often requiring multi-image alignment. Our evaluation across 11 state-of-the-art LVLMs demonstrates a substantial gap between current performance and the proficiency required for robust map understanding. Analysis reveals that these failures extend beyond issues observed in prior VQA datasets, highlighting the need for novel architectures and effective training methods that incorporate cartographic priors and explicit reasoning over map elements. We will release the error taxonomy and baseline results, alongside FRIEDA, to catalyze progress. We encourage the community to build on FRIEDA with methods that explicitly integrate text, symbology, and geospatial structure, toward LVLMs that reason reliably over real-world maps.
@@
Upon the end of the anonymity period, we plan to release: (1) the benchmark (images, QA JSON, taxonomy, and provenance), (2) code for data loading, inference, evaluation, and table/figure generation, (3) code to replicate the annotation interface, and (4) all prompts and configuration files used for annotation and inference. In the meantime, we provide all details needed to reproduce our results in the main text and appendices: Section 3 describes dataset construction, Section 4 specifies models and inference settings, and Section 5 reports ablations and error analyses.
@@ Appendix A Datacard
We adopt the data statement framework of c:bender2018datacard and integrate complementary fields from Datasheets for Datasets (c:gebru2021datasheet) to centralize key information for the analysis, reuse, and deployment of FRIEDA.
@@ B.1 Map Image Filtering
We use Idefics3-8B (c:laurencon2024idefics) to filter map images from the document. To produce a strict Yes/No decision, we prompt the model:
@@ C.3 Nation and Domain Coverages
We source maps from domains where spatial reasoning is essential: geologic reports (c:ni43101), national park management reports (c:national-parks), investment and infrastructure reports (c:aiib), disaster and hazard assessments (c:fema), city and regional planning documents (c:seattle-planning; c:capetown; c:abu-dhabi; c:sg-ura), and environmental reviews (c:eis; c:seychelles; c:ireland). Several of these are umbrella categories that can be further subdivided. For example, environmental assessments may target facilities, hydrology, land use/land cover, or habitat. For consistency, we retain the top-level labels used by the source repositories. Across these domains, maps employ varied symbol conventions (legends, scale bars, north arrows) and heterogeneous geometry types (areas, lines, points), encouraging generalization beyond any single map style. Figure 14 summarizes the domain coverage.
@@ E.1 LLM-as-Judge Prompt
To evaluate free-form textual answers, we employ LLM-as-Judge (c:gu2025surveyllmasajudge) using Mistral-Small-3.1 (c:mistral2024). Since not all models follow our requested output format (“Final answer: ⟨your answer⟩”) and minor wording differences may occur (e.g., ‘15.00%’ vs. ‘15’), we first require the LLM to extract the answer span based on the question and then compare the extracted portion to the gold answer with tolerance for minor variants (Figure 17).

### v2

@@ 4.1 Experimental Setup
We evaluate 11 LVLMs with multi-image support on FRIEDA. For proprietary models, we test three models: Gemini-2.5-Pro (Gemini Team, 2025), GPT-5-Think (OpenAI, 2025), and Claude-Sonnet-4 (Anthropic, 2025). For open source models, we consider eight model families and evaluate the largest available model from each family: LLaVA-NeXT-110B (Li et al., 2024b), GLM4.5V-108B (Team et al., 2025), InternVL3-78B (Chen et al., 2024b), LLaVA-OneVision-72B (Li et al., 2024a), Qwen2.5VL-72B (Bai et al., 2025), InternVL3.5-38B (Wang et al., 2025b), [^2] We evaluate the 38B variant rather than the 241BA28B variant as the latter activates only 28B parameters during inference. We report the results for the 241BA28B setting in Appendix F.2. Ovis2-34B (Lu et al., 2024b), and Ovis2.5-9B (Lu et al., 2025).
To enforce determinism in open-source models, we set do_sample=False and temperature=0. For proprietary models, we use the default settings for each model with maximum reasoning enabled (e.g., reasoning=high for GPT-5-Think) and append the instruction “Do not use search” to disable external retrieval. All models receive the same set of instructions that human annotators receive (Appendix B.4).
@@
Answers in FRIEDA fall into three categories: textual, distance, and direction. For textual answers, we employ an LLM-as-Judge (Gu et al., 2025) method, utilizing Mistral Small 3.1 (Mistral AI, 2024) as the evaluator. [^3] Mistral is not the language backbone of any tested LVLM, thereby reducing potential bias ( Panickssery et al. (2024)). The full judge prompt appears in Appendix E.1. This setup handles minor variation (e.g., ‘Cypress Creek’ vs. ‘Cypress’) by matching semantics rather than identifying exact string equality. For distance-based answer, we report mean absolute percentage error (MAPE) and consider predictions within 20% error as correct, following Lewis (1982). For directional answers, we mark a response correct if it matches the target cardinal direction within one adjacent label (e.g., if the gold answer is North, accept North, North West, and North East), reflecting the perceptual nature of the labels. We validate the reliability of the evaluation method against manual annotations, achieving a Cohen’s $\kappa$ of 0.9028 across all judged questions, which supports its suitability for evaluation.
@@ 4.2 Evaluation Results
Figure 3 summarizes the overall performance, and Table 2 reports accuracy by spatial relation. As FRIEDA retains questions with at least 2/3 annotator agreeing on the gold answer, we report accuracy for two subsets: All-Agree, where all three annotators agreed, and Partial-Agree, where 2/3 annotators agreed. All-Agree items serve as an indirect indicator of questions that are easier and less ambiguous for the annotators under our task and instructions, whereas Partial-Agree items may be considered as intrinsically more difficult or ambiguous to answer correctly. We also report the Overall Accuracy, which aggregates both subsets. Even the strongest LVLM (Gemini-2.5-Pro) remains below 40% overall accuracy, well behind human performance at 84%. The best open-source result (Ovis2.5-9B-Think) achieves 24% overall accuracy, underperforming proprietary systems and falling far below human performance. We find no clear relationship between model size and performance, suggesting that training data, training objectives, and explicit reasoning mechanisms matter more than scale for cartographic reasoning. We present further analyses in the following section and in Appendix F.2.
@@ 5 Analysis
To pinpoint where LVLMs fail, we analyze Gemini-2.5-Pro on the All-Agree subset (167 questions in total). This ensures that our analysis targets distinct model failures on questions that humans find straightforward. We assign each incorrect answer to a single primary error category. When multiple issues co-occur, we prioritize errors that occur earlier in the reasoning pipeline that propagate to downstream steps. The largest source of error is the misinterpretation of legends (25.61%): cases in which the model assigns colors or symbols to the wrong class. The remaining 23.78% is due to cross-map interpretation failures, which reflect difficulties in aligning the map scales and shared features across maps, and 16.46% is due to spatial-relation semantics error, which arises when the model mixes up spatial relations (e.g., labeling region B within A when it only touches A at the boundary). Map-element misunderstandings include mistakes with the map scale (9.76%; unit or ratio errors), map text (8.93%; selecting the wrong place or feature from labels), geometry or shape reference (3.66%; pointing to the wrong area on the map), and orientation (3.05%; ignoring a tilted compass). Finally, we observe generic VQA errors not specific to cartography, such as miscounting (6.71%), subject-object confusion (1.82%; referring ‘A relative to B’ as ‘B relative to A’), and hallucination (1.20%). For the top three error categories, we provide examples and rationales returned by the three proprietary models in Appendix F.1.
Figure 4: Per spatial relation accuracy (%) of human annotators and three proprietary LVLMs (Gemini-2.5-Pro, GPT-5-Think, and Claude-Sonnet-4) on FRIEDA-direct.
@@
Figure 4 reports per-spatial relation accuracy for human annotators and the three proprietary models. LVLM performance broadly tracks the human baseline: both are most accurate on orientation and struggle most with distance. On questions where an annotator answers incorrectly, LVLMs are also incorrect 84.53% of the time. While GPT-5-Think and Gemini-2.5-Pro achieve comparable overall accuracy, GPT-5-Think is stronger on tasks that require multi-map reasoning (Table 10), indicating better integration of evidence across maps. This is most evident in the equal relation questions, a multi-map exclusive task, where GPT-5-Think’s accuracy is nearly 13% higher compared to Gemini-2.5-Pro. Notably, Claude-Sonnet-4 is the strongest on distance questions, particularly those that require interpreting the map scale to compute exact distances.
@@
We observe a minimal difference in accuracy between the FRIEDA-direct (Figure 3 and Table 2) and FRIEDA-contextual (Appendix E.3). To verify that this is not an artifact of the accuracy metric, we directly compare the per-question performance of the eight open-source models under deterministic settings (i.e., do_sample=False and temperature=0). We observe 88.03% agreement in per-question performance between the direct and contextual settings, indicating that contextual images (maps from the same document that are not required to answer the question) rarely affect the model’s predictions.
@@
Despite being the smallest model tested with FRIEDA, Ovis2.5-9B-Think attains strong results (4th overall and 1st among open source models). To identify what drives this performance, we further evaluate Ovis2.5-9B with explicit reasoning (i.e., Think) disabled (Table 3). The overall accuracy of Ovis2.5-9B remains above the open-source average, indicating that model characteristics (e.g., architecture, training data) contribute to its strong performance. Enabling Think adds an additional 5% performance gain. To identify which question types benefit from explicit reasoning and whether it improves cartographic performance, we manually analyze the 60 questions that only the Think variant answers correctly. Reasoning helps mostly with cardinal-direction questions, where north faces the top of the image (48.33%), followed by multi-map alignment (23.33%). Additional improvements arise from correctly reading map text (15%), interpreting the map scale (5%), associating the legend with the symbol (5%), and counting (3%). Together, these patterns suggest that explicit reasoning primarily strengthens orientation-related and multi-map questions, which are central to cartographic reasoning, whereas yielding smaller gains in symbol and map-scale interpretation. [^4] We further evaluate the association between performance and model size in Appendix F.2
@@ 6 Related Work
Recent benchmarks established baselines for LVLM reasoning over documents and designed graphics. In the document domain, DocVQA (Mathew et al., 2021) and DocVXQA (Souibgui et al., 2025) introduce a large-scale question-answering (QA) dataset over real forms and reports, while DocoPilot (Duan et al., 2025) extends evaluation to scientific articles, which involve embedded figures. For graphics, InfographicsVQA (Mathew et al., 2022) and InfoChartQA (Lin et al., 2025) test reasoning over rich layouts and charts. In general, frontier LVLMs reveal competence at high-level patterns, such as trends and extrema, but struggle with precise value extraction and robustness. FRIEDA evaluates these shortcomings in a cartographic setting where layout, symbols, legends, scales, and compass orientation interact tightly to measure how well LVLMs integrate these signals to answer map-based questions.
@@
While recent map VQA benchmarks have expanded the scope of evaluation, they remain constrained to single-map tasks or specific domains. MapQA (Chang et al., 2022) evaluates question answering on choropleth maps and shows that general VQA and ChartVQA systems underperform on maps. MapWise (Mukhopadhyay et al., 2025) broadens the geographic coverage, while MapIQ (Srivastava et al., 2025) extends the map type coverage to include cartograms and proportional-symbol maps. MapEval (Dihan et al., 2025) assesses geospatial reasoning across various cities, and it reports consistent human-LVLMs performance gaps. Domain-specific efforts include PEACE (Huang et al., 2025) for geology, CartoMark (Zhou et al., 2024) for text extraction and recognition, and MapBench (Xing et al., 2025) and ReasonMap (Feng et al., 2025) for navigation. However, these benchmarks rarely test cross-image reasoning on heterogeneous sources and often rely on a limited set of spatial relations. While ReMI (Kazemi et al., 2025) explores the cross-image setting, the questions lack cartographic focus. We detail key differences between prior map VQA benchmarks and FRIEDA in Appendix G.
Spatial reasoning benchmarks, such as SpatialVLM (Chen et al., 2024a) and SpatialRGPT (Cheng et al., 2024), have advanced spatial perception and reasoning on natural images. In the geospatial domain, GeoChain (Yerramilli et al., 2025) enhances tasks like geolocalization by inducing step-by-step geographic reasoning. However, these works do not engage with symbolic conventions unique to maps (i.e., legends, scales, compasses, and map texts). In contrast, our benchmark closes this gap by evaluating multi-step cartographic reasoning over heterogeneous, real-document maps, which requires models to integrate evidence across multiple figures and align legends, scales, and orientation to infer key spatial relations (i.e., border, distance, equal, intersect, orientation, and within).
@@ 7 Conclusion
We present FRIEDA, a benchmark for evaluating multi-step cartographic reasoning over real-world maps across six core spatial relations, where answering often requires aligning evidence across multiple maps. Our evaluation of 11 state-of-the-art LVLMs reveals a substantial gap between current performance and the proficiency required for reliable map understanding. Models frequently fail to ground symbologic map entities, to integrate evidence distributed across multiple maps, and to correctly interpret spatial topology. Our analysis further shows that these failures are not merely amplified versions of errors observed in prior VQA benchmarks, but also expose additional map-specific challenges, including symbol-to-semantics mapping, adherence to cartographic conventions, and geometry-aware reasoning. These findings highlight the need for novel architectures and effective training strategies that incorporate cartographic priors and enable explicit reasoning over map elements and spatial structure. To catalyze progress, we release FRIEDA along with the question taxonomy, baseline results, and evaluation scripts to enable fine-grained diagnosis and reproducible comparisons. We encourage the community to build on FRIEDA with methods that explicitly integrate text, symbology, and geospatial structure, moving toward LVLMs that reason robustly and reliably over real-world maps.
@@
We released the benchmark dataset (images, QA JSON, and taxonomically classified questions) on HuggingFace, along with all prompt and configuration files used for annotation and inference. We also provide the full codebase for data loading, inference, and evaluation on GitHub, as well as code to replicate the annotation interface. The results obtained during testing are released at result branch.
Acknowledgements
This material is based upon works supported in part by the National Science Foundation under Grant No. BCS-2419334.
@@ Appendix A Datacard
We adopt the data statement framework of Bender and Friedman (2018) and integrate complementary fields from Datasheets for Datasets (Gebru et al., 2021) to centralize key information for the analysis, reuse, and deployment of FRIEDA.
@@ B.1 Map Image Filtering
We use Idefics3-8B (Laurençon et al., 2024) to filter map images from the document. To produce a strict Yes/No decision, we prompt the model:
@@ C.3 Nation and Domain Coverages
We source maps from domains where spatial reasoning is essential: geologic reports (SEDAR+, 2025), national park management reports (National Park Service, 2025), investment and infrastructure reports (AIIB, 2025), disaster and hazard assessments (Federal Emergency Management Agency, 2025), city and regional planning documents (Seattle Planning and Community Development, 2024; City of Cape Town, 2025; Department of Community Development, 2025; Urban Redevelopment Authority, 2025), and environmental reviews (U.S. Environmental Protection Agency, 2025; Ministry of Agriculture Climate Change and Environment, 2025; Environmental Protection Agency, 2025). Several of these are umbrella categories that can be further subdivided. For example, environmental assessments may target facilities, hydrology, land use/land cover, or habitat. For consistency, we retain the top-level labels used by the source repositories. Across these domains, maps employ varied symbol conventions (legends, scale bars, north arrows) and heterogeneous geometry types (areas, lines, points), encouraging generalization beyond any single map style. Figure 14 summarizes the domain coverage.
@@ E.1 LLM-as-Judge Prompt
To evaluate free-form textual answers, we employ LLM-as-Judge (Gu et al., 2025) using Mistral-Small-3.1 (Mistral AI, 2024). Since not all models follow our requested output format (“Final answer: ⟨your answer⟩”) and minor wording differences may occur (e.g., ‘15.00%’ vs. ‘15’), we first require the LLM to extract the answer span based on the question and then compare the extracted portion to the gold answer with tolerance for minor variants (Figure 17).